OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (3, 7).

OpenStudy (anonymous):

linear equations like this are easier first get your data pm=2 p(3:7) can use either 1 of the following equations y=mx+c/q or y-y1=m(x-x1) y-7=2{x-3) y-7=2x-6 y=2x-6+7 y=2x+1 or y=mx+c or y=mx+q 7=2(3)+c 7-6=c c=1 thereofer y=2x+1